OpenStudy (anonymous):

suppose one geometric mean g and two arithmetic means l and m are inserted b/n two numbers, find\[g^{2}\] in terms of l and m.

hartnn (hartnn):

m getting g = 2lm/(l+m) g as HM between l and m

OpenStudy (anonymous):

nope

hartnn (hartnn):

x,l,g,m,y g=\(\sqrt{xy}\) l=(x+g)/2 m=(g+y)/2 solving for g, i got g = 2lm/(l+m)

OpenStudy (anonymous):

g^2 =a*b where a and b are the two numbers now is l-(m-l)=2l-m and b is m+m-l =2m-l hence g^2= (2l-m)(2m-l)

OpenStudy (anonymous):

Nope hartnn

OpenStudy (anonymous):

is my answer correct @Zekarias

OpenStudy (anonymous):

That is right matricked

OpenStudy (anonymous):

closing